OpenStudy (anonymous):

For all nonzero real numbers p, t, x, and y such that , (x/y)=(3p/2t) which expression is equivalent to t ?

OpenStudy (anonymous):

Is this a multiple choice question?

OpenStudy (anonymous):

Yes, the choices are: F. y/2 G. 3px/2y H. 6py/x J. 3py/x K. 3py/2x

OpenStudy (anonymous):

Ok well basicall all they're asking for is to solve for t, do you know how to manipluate variables? I can alwsy teach you it's really easy.

OpenStudy (anonymous):

Oh and \[\Huge WELCOME~TO~~~OPENSTUDY\]

OpenStudy (anonymous):

Anyways we have this equation: \[\Large \frac{x}{y}=\frac{3p}{2t}\] correct?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Ok sorry back, now we multiply both sides by 2t to get it from the denominator \[\Huge 2t \times \frac{x}{y}=\frac{3p}{2t} \times 2t\]

OpenStudy (anonymous):

Then we have \[\Huge \frac{2tx}{x}=3p\] then we muliply both sides by \[\Huge \frac{1}{3p} \times \frac{2tx}{x}=3p \times \frac{1}{3p}\] \[\Huge \frac{2tx}{3px}=1\] Then we multiply by 3px/2x to get the "t" by itself \[\Huge \frac{3px}{2x} \times \frac{2tx}{3px}=1 \times \frac{3px}{2x}\] \[\Huge t=\frac{3px}{2x}\]

OpenStudy (anonymous):

Sorry its 3py**** I changed the y into an x

OpenStudy (anonymous):

Letter K

OpenStudy (anonymous):

Thank you.

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."